OpenStudy (anonymous):

In an examination, a student who got 20% of the total marks fails by 18 marks. Had he got 32% of the total marks he would have got 6 marks more than passing marks. What is the passing marks?

OpenStudy (anonymous):

passing marks is 34.8

myininaya (myininaya):

@karthik_a.k.a_joker Please explain your answers. Don't just give an answer without explanation. Try to engage asker. Thanks.

OpenStudy (anonymous):

Equations formed are : P - 0.2x = 18 -P + 0.32x = 6 Adding them, we get the value of x 0.12x = 24 and x = 200 Therefore, Passing marks (P) = 58...

OpenStudy (anonymous):

I mean to say that total marks are 200 and Passing marks are 58..

OpenStudy (anonymous):

let total marks be "x", passing marks be "y" given 20% of total marks means he is failed and it is equal to 18 so it makes (20)(x)/100=18 so x is 90 i.e total marks is 90 given 32% of total marks is equal to 6 marks plus the passing marks so (32)(x)/100=(y+6) we know the value of x get the value if y which is the passing marks

OpenStudy (anonymous):

And one can check the answer too.. 20/100 of 200 = 40 and the passing marks are 58 so he got failed by 18 marks.. 32/100 of 200 = 64, so he would have got 6 marks more...

OpenStudy (anonymous):

@karthik_a.k.a_joker, please check your answer...

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

That is really nice...

OpenStudy (anonymous):

i didnt get you..!! what is nice?